Physics: Given the two charges...

Given the two charges +Q and Q1 = -2Q/3 shown in (
Answer #1

Enet = E1 + E2 = 0

E1 = -E2


Q / (d+x)^2 = 2Q/3*x^2


1/(d+x)^2 = 2/3*x^2

3x^2 = 2d^2 + 2x^2 + 4dx

x^2 - 2d^2 - 4dx = 0


x = (2+sqrt(6))*d


x/d = 2+sqrt(6) = 4.45